OpenStudy (anonymous):

The circuit parameters have the following values: VS=4V and R=1kΩ, and the ExpFET parameters are: K=1mA, VT=1V and VN=1V. Given that vIN=0.5V, what is the value of vOUT in Volts?

OpenStudy (anonymous):

please post the figure....!

OpenStudy (anonymous):

q2 p1: Vs as vin<vt p2: v0 = vs-kr(e^((vin-vt)/vn)-1) p3: vin = ln(vs/2kr +1)vn+vt p4: vin=sqrt(vs/kr)+vt p5: -kre^((vin-vt)/vn) p6: -kr(vin-vt) vs=4.5 r = 1k k =1mA vt=1 vn=1 1: 4.5 vin=0.5 2: 2.7818 vin=2 3: 2.1786 vo=vs/2 4: 3.1213 vo=vs/2 mosfet 5: -2.7182 vin=2 6: -1 vin=2 mosfet Just change the values with yours

OpenStudy (anonymous):

all formullas posted...... http://openstudy.com/study?source=email#/updates/50d4a24de4b0d6c1d5417e42